OpenStudy (anonymous):

Please Help: Decomposition of Functions: f(x)=ABS(x^2-5x)+3 Find g(x) and h(x) such that f(x)=g(h(x)) and neither g(x) or h(x) is solely x

OpenStudy (rogue):

Well, it looks like the inner function, h(x) is...\[h(x) = x^2 - 5x + 3\]The outer function then is,\[g(x) = \left| x \right| + 3\]

OpenStudy (rogue):

oops, \[h(x) = x^2 - 5x\]